OpenStudy (flowey15):

Which point does not lie on the circle centered at (3, 2) with radius 5? A) (0, 6) B) (3, -3) C) (-2, 2) D) (3, 0)

OpenStudy (dinamix):

draw it and u can understand it it's take u little think

OpenStudy (anonymous):

B) (3, -3)

OpenStudy (dinamix):

dont give her direct answer @EpicGammer

OpenStudy (flowey15):

yaay, im a big girl, i got b, lol, jk

OpenStudy (dinamix):

ok let's i check it

OpenStudy (flowey15):

i don't know how to do this :(

OpenStudy (dinamix):

\[(x-3)^2+(y-2)^2>25\] put point and remplce it in this equation if find it greater then 25 its not in circle that's only @flowey15

OpenStudy (dinamix):

u understand or no ? @flowey15

OpenStudy (flowey15):

sorry, i had to find my pen

OpenStudy (dinamix):

@ uare wrong its not B

OpenStudy (dinamix):

i know the answer but it's good when u find it u alone @flowey15

OpenStudy (flowey15):

otay

OpenStudy (dinamix):

A) (0-3)^2+(6-2)^2 = 9+16 =25 that's equal 25 B)(3-3)^2+(-3-2)^2 = 0+25=25 same c) (-2-3)^2 +( 2-2)^2 = 25 +0 same d) (3-3)^2 +(-2)^2 = 0+4 its 4<25